# Ergotropic advantage in a measurement-fueled quantum heat engine

### v1

@@ IV Other measurement directions
We have also studied the performance of the engine with generalized measurements along x-x (y-y), x-y, and x-z (y-z) directions. Usually, for these cases, there are many possibilities for ordering the post-measurement probabilities. To illustrate, we consider x-x measurements for which we set $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{x}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{x}^{\rm B}$ in Eq. (10).The post-measurement probabilities are given by
@@
The relative ordering of the above probabilities is not straightforward, though we can prove specific inequalities, such as $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{3}$. For $c_{0}<1/2$, $p^{\rm{PM}}_{2}<p^{\rm{PM}}_{4}$ ($p^{\rm{PM}}_{2}>p^{\rm{PM}}_{4}$ for $c_{0}>1/2$). For the purpose of illustration, we consider one of the orderings, given by $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{4}\geq p^{\rm{PM}}_{2}>p^{\rm{PM}}_{3}$, in the range $0<c_{0}\leq 1/2$. Firstly, unlike the case of z-z measurements, we have a non-zero total work for the four-stroke cycle with weak measurements. Heat and work involved are shown in Fig. 6. In Fig. 7, we show the effect of weak measurements as they can lead to a higher efficiency compared to projective measurements. The efficiency is plotted versus $J$ (Fig. 8) for a fixed $c_{0}$ to show the effect of the reservoir temperature and we note that the efficiency reduces with the lowering of the temperature.
Figure 6: x-x case: $Q_{\rm M}$ (green curve), $W_{\rm erg}$ (brown curve), $W_{\rm T}$ (red curve), $W_{1}$ (orange curve), $W_{2}$ (black curve) and $Q_{\rm res}$ (blue curve). $B_{1}=3.5$, $B_{2}=3$, $J=1$ and $\beta=1$.
In the special case of x-x projective measurements ($c_{0}=c_{1}=1/2$), Eqs. (24) simplify as
[…]
The heat exchange during the measurement stroke is
@@
The above probability distribution follows a definite ordering: $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{2}=p^{\rm{PM}}_{4}>p^{\rm{PM}}_{3}$, and indicates an active state. The ergotropy is given by $W_{\rm erg}=-B_{1}(p_{1}-p_{3})/2<0$, while the total work output is
$W_{\rm T}^{(5)}=2(B_{1}-B_{2})(p_{2}-p_{4})+\frac{B_{2}}{2}(p_{1}-p_{3})>0.$ (27)
Compared to the above, the work output of a four-stroke cycle with projective measurements is while the three-stroke cycle yields $W_{\rm T}^{(3)}={B_{2}}(p_{1}-p_{3})/2$. Thus, we verify the equality (23).
Similarly, for the x-z projective measurements, we have a definite ordering i.e. $p^{\rm{PM}}_{2}>p^{\rm{PM}}_{1}=p^{\rm{PM}}_{3}(=1/4)>p^{\rm{PM}}_{4}$ which describes an active state. Now, the heat exchange during the measurement stroke is
$Q_{\rm M}=B_{1}(p_{2}-p_{4})+2J(3p_{1}-p_{2}-p_{3}-p_{4}),$ (28)
while the total work output is
$W_{\rm T}^{(5)}=(B_{1}-B_{2})(p_{2}-p_{4})+\frac{1}{2}{(4J-B_{2})}(p_{2}-p_{4}).$ (29)
The first and the second terms on the rhs above can be identified with $W_{\rm T}^{(4)}$ and $W_{\rm T}^{(3)}$ respectively. Thus, we conclude that the projective measurements yield a higher work as well as efficiency for the five-stroke cycle than the corresponding four-stroke cycle.
For the case of x-y projective measurements, we find that the post-measurement probability distribution is uniform ($p_{n}^{\rm PM}=1/4$). Being a passive state, the ergotropic stroke does not help in this case.
Before closing this section, we remark on the case of spin-measurements along arbitrary directions. As mentioned earlier, relations between post-measurement probabilities are hard to obtain here. However, if we restrict ourselves to the projective measurements, the following inequality always holds.
@@
Figure 7: x-x case: Efficiency vs. $J$ for five-stroke engine. Red curve ($c_{0}$ = 0.1), blue curve ($c_{0}=0.3$), black curve ($c_{0}=0.5$), brown curve ($c_{0}=0.6$), green curve ($c_{0}=0.7$). For comparison, the efficiency of a four-stroke engine with projective measurement of Ref. [33] (dashed curve) is also plotted. Here, $B_{1}=3.5$, $B_{2}=3$, and $\beta=1$.
Figure 8: x-x case: Efficiency vs. $J$. Red curve ($\beta=1$), green curve ($\beta=2$), blue curve ($\beta=3$) and magenta curve ($\beta=4$). $B_{1}=3.5$, $B_{2}=3$ and $c_{0}=0.3$.
V Summary
In this paper, we have studied a measurement-fueled quantum engine based on a coupled two-qubits system where weak measurements of the spin components are used as a source of heat. In the specific cases we have analyzed, the z–z case yields zero work output in the four-stroke cycle with generalized measurements. There are two possible orderings of the post-measurement probabilities which correspond to active post-measurement state $\rho^{PM}$. Incorporating an ergotropy-extracting stroke after the measurement stroke yields a positive work output. Furthermore, a reduced three-stroke cycle (without adiabatic strokes) yields the same performance as the five-stroke cycle.
For measurement directions, such as x–x or x–z, the four-stroke cycle produces non-zero work, and the ergotropic extension further improves the performance. In general, several orderings of the post-measurement probabilities may occur, although projective measurements can yield a unique ordering. Notably, weak x–x and x–z measurements within a five-stroke cycle can achieve a higher efficiency than projective measurements, underscoring the role of weak measurements in optimizing the performance of a quantum engine.

### v2

@@ IV Other measurement directions
We also studied the performance of the engine with generalized measurements along x-x (y-y), x-y, and x-z (y-z) directions. In general, the post-measurement state $\rho^{\rm PM}$ contains some coherence (in the energy basis), making it an active state. Then, to determine the passive state after the ergotropic stroke, we need to order the eigenvalues of the state $\rho^{\rm PM}$ in a non-increasing sense. To illustrate, we consider x-x measurements for which we set $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{x}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{x}^{\rm B}$ in Eq. (10). For the case of x-x weak measurements, the state $\rho^{\rm PM}$ is active, and we get the post-measurement occupation probabilities as
@@
The relative ordering of the above probabilities is not straightforward, though we can prove specific inequalities, such as $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{3}$. For $c_{0}<1/2$, $p^{\rm{PM}}_{2}<p^{\rm{PM}}_{4}$ while $p^{\rm{PM}}_{2}>p^{\rm{PM}}_{4}$ for $c_{0}>1/2$. Unlike the case of z-z measurements, here the four-stroke cycle with weak measurements yields a non-zero total work.
The eigenvalues of $\rho^{\rm PM}$ can be written as
$\begin{split}p_{1}^{\prime}&=p^{\rm{PM}}_{1},\\ p_{2}^{\prime}&=\frac{p^{\rm{PM}}_{2}+p^{\rm{PM}}_{4}}{2}+\Delta,\\ p_{3}^{\prime}&=p^{\rm{PM}}_{3},\\ p_{4}^{\prime}&=\frac{p^{\rm{PM}}_{2}+p^{\rm{PM}}_{4}}{2}-\Delta,\end{split}$ (26)
where $\Delta>0$ depends in a complicated way on $p_{n}$ and $c_{0}$. The above eigenvalues are also the occupation probabilities in the passive state $\rho^{\prime}$, to be ordered in a decreasing sense. However, there can be different orderings of these probabilities in different parameter regimes. On the other hand, for the specific case of projective measurements ($c_{0}=1/2$), the post-measurement probabilities are simplified as: $p^{\rm{PM}}_{1}=(1+p_{1}-p_{3})/4,\;p^{\rm{PM}}_{2}=p^{\rm{PM}}_{4}=1/4,\;p^{\rm{PM}}_{3}=(1-p_{1}+p_{3})/4$, while the eigenvalues of $\rho^{\rm PM}$ are $p_{1}^{\prime}=p_{2}^{\prime}=p^{\rm{PM}}_{1}$ and $p_{3}^{\prime}=p_{4}^{\prime}=p^{\rm{PM}}_{3}$, which are already ordered i.e. $p_{1}^{\prime}=p_{2}^{\prime}>p_{4}^{\prime}=p_{3}^{\prime}$, since $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{3}$. In this special case, the heat exchanged during the measurement stroke is calculated to be
@@
The ergotropy is given by $W_{\rm erg}=-B_{1}(p_{1}-p_{3})<0$, while the total work output is
$W_{\rm T}^{(5)}=2(B_{1}-B_{2})(p_{2}-p_{4})+B_{2}(p_{1}-p_{3})>0.$ (28)
The work output of a four-stroke cycle with projective measurements is $W_{\rm T}^{(4)}=2(B_{1}-B_{2})(p_{2}-p_{4})<W_{\rm T}^{(5)}$, while the three-stroke cycle yields $W_{\rm T}^{(3)}={B_{2}}(p_{1}-p_{3})$. Thus, we verify the equality (24).
In Fig. 6, we show the effect of weak measurements as they can lead to a higher efficiency compared to projective measurements. Here, the ordering of the probabilities in the post-ergotropy state is performed numerically for the given choice of system parameters. Similarly, we find that the post-measurement state for the x-z case has coherence and so is an active state. On the other hand, with x-y projective measurements, we find that the post-measurement probability distribution is uniform ($p_{n}^{\rm PM}=1/4$). Being a passive state, the ergotropic stroke does not help in this case.
Before closing this section, we consider the case of spin-measurements along arbitrary directions. As mentioned earlier, ordering relations between the post-measurement probabilities or between the eigenvalues of $\rho^{\rm PM}$ are hard to obtain here. However, if we restrict ourselves to projective measurements, the following inequality always holds.
@@
Figure 6: x-x case: Efficiency vs. $J$ for the five-stroke engine. Weak measurements with $c_{0}<1/2$ can lead to a higher efficiency than with projective measurements ($c_{0}=1/2$). For comparison, the efficiency of the four-stroke engine with projective measurements, $\eta^{(4)}$, is also plotted. Here, $B_{1}=3.5$, $B_{2}=3$ and $\beta=1$.
V Conclusions
In this paper, we studied the effect of additional ergotropy extraction on the performance of a measurement-fueled quantum engine based on a coupled two-qubits system. Here, the weak measurements of the spin components are used as a resource for input heat. We first analyzed the case of z-z weak measurements in a five-stroke cycle. The corresponding four-stroke cycle yields zero work output. Notably, the post-measurement state has no coherence in the energy eigenbasis. However, there are two possible orderings of the post-measurement probabilities leading to an active post-measurement state. We find that an ergotropy-extracting stroke inserted after the measurement stroke enables a net positive work output. Furthermore, a reduced three-stroke cycle (without the adiabatic strokes) yields the same performance as the five-stroke cycle. However, for other measurement directions (e.g., x–x, x–z), the post-measurement active state, in general, has coherence. Here, the four-stroke engine already delivers non-zero work, yet the inclusion of an ergotropic stroke enhances the performance. We also find that weak x–x measurements can surpass the projective ones in efficiency, highlighting the role of the measurement strength as a tuning parameter. In contrast, some settings (e.g., x–y case with projective measurements) yield passive post-measurement states, offering no ergotropic advantage.
A general identity, valid for arbitrary Hamiltonian and non-selective measurements, was established, which shows that the total work of the five-stroke engine equals the sum of its four-stroke and three-stroke counterparts. The present findings underline three key points: (i) measurements combined with ergotropy extraction can increase useful work; (ii) engine performance depends sensitively on the measurement strength; and (iii) appropriate choices of these parameters can outperform both standard four-stroke and three-stroke cycles.
In our analysis, we have neglected some of the energetic and thermodynamic costs towards the implementation of the unitary protocols [62, 63]. The notion of work extraction from a quantum system assumes some coupling with a work reservoir, or a quantum battery. In certain situations, with a battery of a finite size and a finite coherence, the work extracted is expected be less than ergotropy [64]. One possibility is to assume that the battery is large enough and highly coherent, so that all the energy transferred to the battery can be interpreted as work. Further, the post-measurement state of the working medium may have some degree of coherence. Then, extracting the coherent part of the ergotropy also becomes an experimental challenge. We have assumed a quasi-static heat cycle. Thus, in the present framework, we take the costs incurred through these time-dependent protocols [65] to be negligible. However, in the corresponding finite-time thermodynamic cycle, these costs can be important, and accordingly, the comparison between the performance of four-stroke and five-stroke engines has to be reassessed.
In conclusion, we have demonstrated that combining measurement-induced population reshaping with the extraction of ergotropy provides a viable thermodynamic advantage for quantum engines. These features can be exploited in feasible experimental platforms such as NMR [51] and superconducting circuits [65] which have so far formulated state of the art protocols for single qubit cases. Future work can include many-body systems, feedback-based protocols and bounding the performance due to finite-time effects.